# Q-Net: Transferable Queue Length Estimation via Kalman-based Neural Networks

### v1

@@ 3 Problem setup and challenges
A lane refers to a portion of the roadway designated for a single line of vehicles, which serves to organize traffic, guide drivers, and reduce conflicts. A section is a group of lanes that share the same direction of travel. As illustrated in Figure 1, a single lane may accommodate multiple movement directions, such as through and right-turn traffic. A section typically contains multiple lanes, which may diverge into different movement directions at the downstream intersection.
@@
Our problem of interest is to estimate queue lengths at signalized intersections using indirect data sources that do not explicitly measure queues. In urban road traffic networks, direct queue measurement is often impractical or costly, yet accurate queue length estimation is essential for traffic management. We focus on section-wise queue length estimation, where the queue length represents the maximum queuing distance across all lanes within a section.
The problem involves fusing heterogeneous data sources with different characteristics. Loop detectors installed near stop lines record vehicle passages. These passages are aggregated into 10-second vehicle counts. The aFCD divides each road section into segments and provides average speed data for each segment at 1-minute intervals. Radar sensors capture ground truth queue lengths at 10-second intervals. A summary is presented in Table 1.
This presents several key challenges from both road network structure and data characteristics:
@@
    Partially observed traffic flow. Figure 1 illustrates scenarios leading to unobserved vehicle movements: (i) vehicles not entering downstream areas due to multi-directional upstream lanes, and (ii) vehicles exiting or entering the study area via neighboring environments that bypass the monitored network. Entering vehicles typically form minor flows, while exiting vehicles contribute to secondary flows. These unmonitored movements introduce uncertainty into queue length estimation.
@@
    Spatiotemporal data inconsistency. The data sources exhibit significant heterogeneity in both spatial granularity and temporal resolution, as shown in Table 1. This inconsistency requires models capable of reconciling spatial and temporal mismatches for effective data fusion.
Our goal is to develop a data-driven framework that effectively combines these multi-resolution, multi-source data to produce accurate queue length estimates. The following section presents our proposed approach to address these challenges, highlighting methodological innovations and practical relevance for real-world applications.
@@ 4 Q-Net state-space formulation
Let $x_{t}\in\mathbb{R}$ be the section-level queue length at time step $t$ as the state variable, let $u_{t}\in\mathbb{R}$ be the queue change as the control input, and let $\mathbf{y}_{t}=[y_{t}^{1},y_{t}^{2},\dots,y_{t}^{N}]^{\top}\in\mathbb{R}^{N}$ be the segment-wise aFCD speeds as measurements, where $N$ is the number of section segments. Our state-space model has the form:
@@
This formulation seamlessly combines spatiotemporally inconsistent data sources within the Q-Net framework. The evolution equation (9) expresses the queue length $x_{t}$ at time $t$ as the sum of the previous queue length $x_{t-1}$, the queue change $u_{t}$, and the process noise $w_{t}$, while the measurement equation (10) establishes the relationship between aFCD speeds $\mathbf{y}_{t}$ and queue length $x_{t}$ through a nonlinear mapping $\mathbf{h}(\cdot)$.
The remaining model components are defined as follows: $w_{t}\sim\mathcal{N}(0,Q_{t})$ is Gaussian process noise with covariance $Q_{t}$, and $\mathbf{v}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{t})$ is Gaussian measurement noise with covariance $\mathbf{R}_{t}$. The control input $u_{t}$ (queue change) is derived from cumulative loop detector counts (see Subsection 4.1), and $\mathbf{h}(\cdot)$ is a nonlinear mapping function detailed in Subsection 4.2.
We do not treat noise covariances as fixed or known; instead, $Q_{t}$ and $\mathbf{R}_{t}$ are treated as time-varying and learnable parameters. The covariance of state variable and measurements are denoted as $\Sigma_{t}$ and $\mathbf{S}_{t}$. A summary of the notations used throughout this paper is provided in Table 2.
@@
$w_{t},Q_{t}$ Process noise and its covariance matrix m, m²
@@
$\Sigma_{t}$ Second-order statistical moment of state variable $x_{t}$
$\mathbf{S}_{t}$ Second-order statistical moment of measurements $\mathbf{y}_{t}$ (m/s)²
@@
$Q_{\max}$ Maximum queue length for the section m
@@ 4.1 Control queue change input
In this subsection, we explain how to derive the control input $\mathbf{u}_{t}$ in Equation (9). The approach involves recovering queue length using only cumulative vehicle counts and utilizing the temporal differences as queue changes. Since we are primarily interested in the control input (i.e., queue change), our goal is not to reproduce absolute queue length values, but rather to capture the temporal dynamics of queue buildup and dissipation. Let $q_{t}$ denote the reconstructed queue length at time $t$. The control input is derived as:
$u_{t}=\mathcal{D}(\mathcal{F}(\mathcal{C}(q_{t}))),$ (11)
where $\mathcal{C}$ is the operation to account for unobserved traffic flow, $\mathcal{F}$ is the Fourier band-pass filtering operation, and $\mathcal{D}$ is the temporal difference operation.
@@
Queue length recovery considering unobserved traffic. Let $\lambda_{c}$ be the unobserved net traffic flow rate, defined as the net number of vehicles leaving the system per unit time without passing through the designated outflow point, partially offset by unobserved minor traffic inflow. We model the section with length $L$ and $m$ lanes with equally formed queues, where the queued portion operates at jam density $k_{j}$ and speed $v_{j}$, while the remaining portion operates at free-flow density $k_{f}$ and speed $v_{f}$. Let $A_{t}$ and $D_{t}$ respectively be cumulative vehicle inflow and outflow up to time $t$. Applying macroscopic vehicle conservation:
$mq_{t}k_{j}+m(L-q_{t})k_{f}=A_{t}-D_{t}-\lambda_{c}t.$ (12)
@@
The lefthand side represents the total amount of vehicles in the section: $mq_{t}k_{j}$ in the queued region and $m(L-q_{t})k_{f}$ in the free-flow region. The righthand side accounts for net vehicle accumulation adjusted for unobserved flows. Equation (12) yields:
$q_{t}=\frac{A_{t}-D_{t}-mLk_{f}-\lambda_{c}t}{m(k_{j}-k_{f})}.$ (13)
@@
Since $A_{t}-D_{t}$ can be directly obtained from loop detector counts, the only remaining unknown that governs the temporal evolution of $q_{t}$ is the unobserved net flow rate $\lambda_{c}$
We estimate $\lambda_{c}$ using boundary conditions where queue length is approximately zero at the beginning and end of each analysis period (e.g., start and end of day). Once $\lambda_{c}$ is identified, rather than fixing the values for $k_{j}$ and $k_{f}$, we apply an affine transformation to $A_{t}-D_{t}-\lambda_{c}t$ through shifting and scaling so that the adjusted signal spans the empirically observed queue length range $0,Q_{\max}]$, where $Q_{\max}$ is the section-specific maximum. This approach avoids assumptions about jam and free-flow densities, allowing the model to learn these parameters directly from daily vehicle count patterns. The reconstructed queue length is shown in Figure 2(a).
@@
(a) Orange line: Raw net vehicle accumulation $(A_{t}-D_{t})$. Blue line: Reconstructed queue length by removing the unobserved traffic flow from the net flow and applying an affine transformation to align with the observed queue range.
@@
(b) Orange line: Reconstructed queue length using a Fourier filter that suppresses low-frequency drift and high-frequency noise, producing a smoothed queue profile suitable for control input $u_{t}$. Blue line: Ground truth queue length.
@@
Control input derivation from reconstructed queue length. From Figure 2(a), the previous reconstructed queue length exhibits a slow-varying trend over time. This low-frequency drift is likely due to the fact that the unobserved traffic flow rate $\lambda_{c}$ is not strictly constant over time—e.g., due to changes in turning ratios at different periods during the day. Additionally, high-frequency noise may arise from detector errors and aggregation mismatches. We apply a Fourier-based bandpass filter to suppress both low-frequency drift and high-frequency noise, emphasizing medium-frequency variations most representative of actual queue dynamics. Figure 2(b) shows the filtered reconstructed queue length. Since our goal is to approximate queue changes (the control input) rather than exact magnitudes, even if the average estimated queue during the afternoon peak appears similar to that in the morning, this is of limited importance because what matters are the relative changes.
@@ 4.2 Measurement function
The equation shows that fully unqueued or queued segments have expected speed observation of $v_{f}$ or $v_{j}$. For partially queued segments, where the queue boundary lies within the segment, the average speed is calculated as segment length divided by total travel time, which is the sum of time in the queued portion (traveled at $v_{j}$) and unqueued portion (traveled at $v_{f}$). Applying this formulation to all segments, we build a mapping function from queue length $x_{t}$ to the measurement vector $\mathbf{{y}}_{t}=[{y}^{1}_{t},{y}^{2}_{t},\dots,{y}^{N}_{t}]^{\top}$.
The characteristic speeds, $v_{j}$ and $v_{f}$, are empirically estimated by picking the two dominant modes from the observed aFCD speed distribution through histogram analysis.
@@ 5 Queue-length Estimation framework
    • $\hat{x}_{t|t-1}=\min(\max(\hat{x}_{t|t-1}+u_{t},0),Q_{\max})$
    • $\hat{\mathbf{y}}_{t|t-1}=\mathbf{h}(\hat{x}_{t|t-1})$
@@
    • $\hat{x}_{t|t}=\hat{x}_{t|t-1}+\mathcal{K}_{t}(\mathbf{y}_{t}-\hat{\mathbf{y}}_{t|t-1})$
    • $\hat{x}_{t|t}=\min(\max(\hat{x}_{t|t},0),Q_{\max})$
@@
$\hat{x}_{t|t-1}=$ $\min(\max(\hat{x}_{t-1|t-1}+u_{t},0),Q_{\max}),$ (15)
$\hat{\mathbf{y}}_{t|t-1}$ $=\mathbf{h}(\hat{x}_{t|t-1}).$ (16)
A clipping operation is applied to ensure that the state variable remains within the range $0,Q_{\max}]$.
Update. As for the update step, the system state is updated using the innovation (i.e., the difference between the predicted and observed measurements) scaled by the Kalman gain $\mathcal{K}_{t}$. Classical Kalman filtering imposes linear evolution and measurement models, along with fixed Gaussian noise. Under these assumptions, the Kalman gain can be computed analytically from the second-order statistics of the state and observation variables. However, in queue length estimation, the state-space model is nonlinear (see Section 4), and the noise characteristics can vary significantly over time. For instance, as shown in Figure 2, the process noise during the afternoon differs significantly from that in the morning. To address these challenges, we learn the Kalman gain $\mathcal{K}_{t}(\theta)$ parameterized by $\theta$ directly from the data following the KalmanNet principle. After receiving the true measurement $\mathbf{y}_{t}$, the system state is updated by:
$\hat{x}_{t|t}=\min(\max(\hat{x}_{t|t-1}+\mathcal{K}_{t}(\theta)(\mathbf{y}_{t}-\mathbf{\hat{y}}_{t|t-1}),0),Q_{\max}).$ (17)
The min-max operation implements state clamping compared to the classical Kalman filter update (see Equation (7)). While this modification sacrifices mathematical optimality, it ensures state estimates remain within physically meaningful bounds $0,Q_{\max}]$. This approach is widely accepted in engineering practice due to its simplicity and computational efficiency (Simon, 2010; Prakash et al., 2014). Since the true measurements (aFCD) have a coarser temporal resolution (60s) compared to the state variable (queue length estimated every 10s), we use the same speed measurement for all time steps within the corresponding one-minute interval.
Within this framework, Q-Net integrates data sources of different resolutions: loop detector data, available at high temporal resolution, is used in the predict step to model queue evolution, while coarse aFCD speed measurements are used to correct predictions in the update step. This design ensures that the overall trend of queue dynamics is shaped by speed observations, while finer temporal variations are captured through detailed loop detector data.
The detailed computation of $\mathcal{K}_{t}(\theta)$ is detailed in Subsection 5.1, and the adaptations to enforce spatial transferability are described in Subsection 5.2.

### v2

@@ 3 Problem setup and challenges
This section presents the problem formulation for section-wise queue length estimation using heterogeneous data sources. We first define the road network terminology, specify the estimation objective, and then describe the data characteristics and associated challenges.
A lane refers to a portion of the roadway designated for a single line of vehicles, which organizes traffic, guides drivers, and reduces conflicts. A section is a group of lanes that share the same direction of travel. As illustrated in Figure 1, a single lane may accommodate multiple movement directions. A section typically contains multiple lanes, which may diverge into different movement directions at the downstream intersection.
@@
Our problem of interest is to estimate queue lengths at signalized intersections using indirect data sources that do not explicitly measure queues. Direct queue measurement is often impractical or costly, yet accurate queue length estimation is essential for urban traffic management. We focus on section-wise queue length estimation, where the queue length represents the maximum queuing distance across all lanes within a section. The problem involves fusing heterogeneous data sources with different characteristics. Loop detectors installed near stop lines record vehicle passages. These passages are aggregated into 10-second vehicle counts. The aFCD divides each road section into segments and provides average speed data for each segment at 1-minute intervals. Radar sensors capture ground truth queue lengths at 10-second intervals. A summary of the data is presented in Table 1.
Queue length estimation under this problem setup is challenging due to limitations in both the road network structure and the data. Broadly, three categories of challenges arise in traffic data fusion: incomplete observability, spatiotemporal heterogeneity, and sensor noise:
@@
    Partial observability of traffic flows. Figure 1 illustrates scenarios leading to unobserved vehicle movements: (i) vehicles not entering downstream areas due to multi-directional upstream lanes, and (ii) vehicles exiting or entering the study area via neighboring environments that bypass the monitored network. Unmonitored entering vehicles typically constitute minor inflows, while unmonitored exiting vehicles can represent substantial outflows. These unmonitored flows propagate uncertainty into queue length estimation and make direct aggregation of detector data unreliable.
@@
    Spatiotemporal heterogeneity of data. The data sources exhibit significant heterogeneity in both spatial granularity and temporal resolution, as shown in Table 1. This inconsistency requires models capable of reconciling spatial and temporal mismatches for effective data fusion.
  3. 3.
    Sensor noise. Each data source is subject to measurement noise: loop detectors may miscount due to calibration errors or occlusion, aFCD may suffer from reporting delays that distort speed estimates, and radar sensors can mismeasure queues when limits are exceeded. Such imperfections complicate data fusion and call for models robust to noisy inputs.
Our goal is to develop a data-driven framework that effectively combines these multi-resolution, multi-source data to produce accurate queue length estimates. To achieve this, we first introduce a novel state-space formulation describing how traffic queue dynamics are represented in Q-Net (Section 4), and then present our AI-augmented Kalman filtering approach for estimating queue length (Section 5).
@@ 4 Q-Net state-space formulation
Let $x_{t}\in\mathbb{R}$ be the section-level queue length at time step $t$ as the state variable, $u_{t}\in\mathbb{R}$ the queue change as the control input, and $\mathbf{y}_{t}=[y_{t}^{1},y_{t}^{2},\dots,y_{t}^{N}]^{\top}\in\mathbb{R}^{N}$ the segment-wise aFCD speeds as measurements over $N$ section segments. Our state-space model has the form:
@@
This formulation combines spatiotemporally inconsistent data sources within the Q-Net framework. The evolution equation (11) expresses the queue length $x_{t}$ at time $t$ as the sum of the previous queue length $x_{t-1}$, the queue change $u_{t}$, and the process noise $w_{t}$, while the measurement equation (12) establishes the relationship between aFCD speeds $\mathbf{y}_{t}$ and queue length $x_{t}$ through a nonlinear mapping $\mathbf{h}(\cdot)$. The remaining model components are $w_{t}\sim\mathcal{N}(0,Q_{t})$ is Gaussian process noise with variance $Q_{t}$, and $\mathbf{v}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{t})$ is Gaussian measurement noise with covariance matrix $\mathbf{R}_{t}$. The control input $u_{t}$ (queue change) is derived from cumulative loop detector counts (see Subsection 4.1), and $\mathbf{h}(\cdot)$ is a nonlinear mapping function detailed in Subsection 4.2.
We do not treat noise covariances as fixed or known; instead, $Q_{t}$ and $\mathbf{R}_{t}$ are time-varying and will be learned from the data. The variance and covariance matrix of state variable and measurements are denoted respectively as $\Sigma_{t}$ and $\mathbf{S}_{t}$. A summary of the notations used throughout this paper is provided in Table 2.
@@
$w_{t},Q_{t}$ Process noise and its variance matrix m, m²
@@
$L^{q}_{t}$ Recovered queue length using only cumulative vehicle counts m
$\Sigma_{t}$ Variance of state variable $x_{t}$
$\mathbf{S}_{t}$ Covariance matrix of measurements $\mathbf{y}_{t}$ (m/s)²
@@
$Q_{\max}$ Maximum reachable queue length for the section m
@@ 4.1 Control queue change input
In this subsection, we explain how to derive the control input ${u}_{t}$ in Equation (11). The approach involves recovering queue length using only cumulative vehicle counts and utilizing the temporal differences as queue changes. While absolute queue lengths and their changes are relevant, our focus is on queue changes, since they determine the control input. Absolute queue lengths are still reconstructed as part of the process, but the temporal dynamics of buildup and dissipation are most critical for control.
Let $A_{t}$ and $D_{t}$ respectively be cumulative vehicle inflow and outflow up to time $t$. Let $L^{q}_{t}$ denote the reconstructed queue length at time $t$. The control input $u_{t}$ is obtained through the following steps:
$L^{q}_{t}$ $=\mathcal{C}(A_{t}-D_{t}),$ (13)
$\bar{L}^{q}_{t}$ $=\mathcal{F}(L^{q}_{t}),$ (14)
$u_{t}$ $=\bar{L}^{q}_{t}-\bar{L}^{q}_{t-1},$ (15)
where $\mathcal{C}$ reconstructs $L^{q}_{t}$ from vehicle count data while accounting for unobserved flows, $\mathcal{F}$ applies a Fourier band-pass filter to produce the smoothed queue length $\bar{L}^{q}_{t}$, and $u_{t}$ is computed as the temporal difference of $\bar{L}^{q}_{t}$. This order ensures corrections and noise suppression precede differencing, yielding a noise-robust measure of queue change that serves as the control input in Equation (11).
@@
Step 1: Queue length recovery $\mathcal{C}(A_{t}-D_{t})$. Let $\lambda_{c}$ be the unobserved net traffic flow rate, defined as the net number of vehicles leaving the system per unit time without passing through the designated outflow point, partially offset by unobserved minor traffic inflow. We model the section of length $L$ with $m$ lanes, considering that queues are approximately equal across lanes. This simplification is reasonable because vehicles can switch lanes downstream for different turning movements, which tends to balance queue lengths over the section. The queued portion operates at jam density $k_{j}$ and speed $v_{j}$, while the remaining portion operates at free-flow density $k_{f}$ and speed $v_{f}$. Applying macroscopic vehicle conservation, we have:
$mL^{q}_{t}k_{j}+m(L-L^{q}_{t})k_{f}=A_{t}-D_{t}-\lambda_{c}t.$ (16)
@@
The left-hand side represents the total amount of vehicles in the section: $mL^{q}_{t}k_{j}$ in the queued region and $m(L-L^{q}_{t})k_{f}$ in the free-flow region. The right-hand side accounts for net vehicle accumulation adjusted for unobserved flows. Equation (16) yields the reconstructed queue length expression:
$L^{q}_{t}=\frac{A_{t}-D_{t}-mLk_{f}-\lambda_{c}t}{m(k_{j}-k_{f})}.$ (17)
@@
Step 1.1: Estimate $\lambda_{c}$. Since $A_{t}-D_{t}$ can be obtained from loop detector counts, the only remaining unknown that governs the temporal evolution of $L^{q}_{t}$ in Equation (17) is the unobserved net flow rate $\lambda_{c}$. We estimate it by applying boundary conditions at the beginning $t_{0}$ (e.g., 6:00 AM) and end $t_{T}$ (e.g., 8:00 PM) of each daily analysis period. This relies on the physical assumption that queues are negligible ($L^{q}_{t}\approx 0$) at these pre-peak ($t_{0}$) and post-peak ($t_{T}$) times. Applying these conditions yields: $\lambda_{c}\approx\frac{(A_{t_{T}}-D_{t_{T}})-(A_{t_{0}}-D_{t_{0}})}{t_{T}-t_{0}}$.
Step 1.2: Estimate $L^{q}_{t}$. Equation (17) conceptually establishes that $L^{q}_{t}$ is an affine transformation of $(A_{t}-D_{t}-\lambda_{c}t)$. Rather than estimating ($k_{j},k_{f}$) in this equation, which can be time-varying and site-specific, we implement this principle directly. The reconstructed queue length $L^{q}_{t}$ is obtained by applying this affine transformation to $(A_{t}-D_{t}-\lambda_{c}t)$, shifting and scaling the signal to span the empirically observed range $0,Q_{\max}]$, where $Q_{\max}$ is the section-specific maximum reachable queue length. The resulting corrected queue length $(L^{q}_{t})$ is shown in Figure 2(a).
Step 2: Fourier band-pass filtration $\mathcal{F}(L^{q}_{t})$. From Figure 2(a), $L^{q}_{t}$ exhibits a slow-varying trend over time. This low-frequency drift is caused by the unobserved traffic flow rate $\lambda_{c}$ varying over time, e.g., due to changes in turning ratios at different periods during the day. Additionally, high-frequency noise may arise from detector errors. We apply a Fourier-based bandpass filter $\mathcal{F}(\cdot)$ to suppress both low-frequency drift and high-frequency noise, emphasizing medium-frequency variations most representative of actual queue dynamics. Figure 2(b) shows the filtered reconstructed queue length $\mathcal{F}(L^{q}_{t})$. Since our goal is to approximate queue changes (the control input) rather than exact magnitudes, even if the average estimated queue during the afternoon peak appears similar to that in the morning, this is of limited importance because what matters are the relative changes.
@@
(a) Reconstructed queue length $L^{q}_{t}$ is obtained by removing the unobserved traffic flow and applying an affine transformation to align with the observed queue range.
@@
(b) Fourier filter suppresses low-frequency drift and high-frequency noise, producing a smoothed queue profile $\bar{L}^{q}_{t}$ suitable for control input $u_{t}$.
@@ 4.2 Measurement function
The equation shows that fully unqueued or queued segments have expected speed observation of $v_{f}$ or $v_{j}$ [^1] The characteristic speeds, $v_{j}$ and $v_{f}$, are empirically estimated by picking the two dominant modes from the observed aFCD speed distribution through histogram analysis.. For partially queued segments, where the queue boundary lies within the segment, the average speed is calculated as segment length divided by total travel time, which is the sum of time in the queued portion (traveled at $v_{j}$) and unqueued portion (traveled at $v_{f}$). Applying this formulation to all segments, we build a mapping function from queue length $x_{t}$ to the measurement vector $\mathbf{{y}}_{t}=[{y}^{1}_{t},{y}^{2}_{t},\dots,{y}^{N}_{t}]^{\top}$.
@@ 5 Queue-length estimation framework
    • $\hat{x}_{t|t-1}=\mathbb{P}_{[0,Q_{\max}]}(\hat{x}_{t-1|t-1}+u_{t})$ $\leftarrow$ Equation (19)
    • $\hat{\mathbf{y}}_{t|t-1}=\mathbf{h}(\hat{x}_{t|t-1})$ $\leftarrow$ Equation (18),(20)
@@
    • $\text{context}_{t}=(\Delta\tilde{x}_{t},\Delta\hat{x}_{t},\Delta\tilde{\mathbf{y}}_{t},\Delta{\mathbf{y}}_{t})$ $\leftarrow$ Equation (22),(23),(24),(25)
@@
    • $\hat{x}_{t|t}=\mathbb{P}_{[0,Q_{\max}]}(\hat{x}_{t|t-1}+\mathcal{K}_{t}(\mathbf{y}_{t}-\hat{\mathbf{y}}_{t|t-1}))$ $\leftarrow$ Equation (21)
@@
$\hat{x}_{t|t-1}$ $=\mathbb{P}_{[0,Q_{\max}]}(\hat{x}_{t-1|t-1}+u_{t}),$ (19)
$\hat{\mathbf{y}}_{t|t-1}$ $=\mathbf{h}(\hat{x}_{t|t-1}),$ (20)
where $\mathbb{P}_{[0,Q_{\max}]}(\cdot)$ is a projection onto the interval $0,Q_{\max}]$, ensuring the state variable remains within its physical range (i.e., the queue length cannot be negative or exceed the section capacity). This ad hoc projection function makes the prior estimate nonlinear, rendering it sub-optimal under classical Kalman filter theory. Nevertheless, the underlying approximation is conceptually similar to other Kalman filter extensions, such as the Extended Kalman Filter, which employs a linearized function to approximate the true dynamics (Maybeck, 1982). Furthermore, while the projection renders our prior estimates nonlinear, we employ a data-driven approach using KalmanNet (Revach et al., 2022), which learns the Kalman gain directly from data. This learning-based framework does not rely on linearity assumptions, and the neural network is expected to adapt to such nonlinear effects during training.
Update. The system state is updated using the innovation (i.e., the difference between the predicted and observed measurements) scaled by the Kalman gain $\mathcal{K}_{t}$. In queue length estimation, the state-space model is nonlinear (see Section 4), and the noise characteristics can vary significantly over time. For instance, as shown in Figure 2, the process noise during the afternoon differs significantly from that in the morning. To address these challenges, we learn the Kalman gain $\mathcal{K}_{t}(\theta)$ parameterized by $\theta$ directly from the data following the KalmanNet principle. After receiving the measurement $\mathbf{y}_{t}$, the system state is updated by:
$\hat{x}_{t|t}=\mathbb{P}_{[0,Q_{\max}]}(\hat{x}_{t|t-1}+\mathcal{K}_{t}(\theta)(\mathbf{y}_{t}-\mathbf{\hat{y}}_{t|t-1})).$ (21)
Similar to the prediction step, the projection on state ensures state estimates remain within physically meaningful bounds, is widely accepted in engineering practice due to its simplicity and computational efficiency (Simon, 2010; Prakash et al., 2014).
Loop detector data are available at a 10s interval, while aFCD measurements arrive at a coarser 60s resolution. To fuse these data, Q-Net performs predict and update steps every 10s. When no new measurement is observed, the most recent aFCD measurement $\mathbf{y}_{t}$ is held and re-used for the update. This provides a meaningful correction because the predicted measurement $\hat{\mathbf{y}}_{t|t-1}$ evolves at each step, resulting in a non-zero innovation ($\mathbf{y}_{t}-\hat{\mathbf{y}}_{t|t-1}$). Although aFCD data are coarse and can be delayed, the high-frequency loop detector inputs used in the prediction step help compensate for this latency. This design enables Q-Net to effectively fuse heterogeneous data sources: high-resolution loop detector data capture rapid queue fluctuations in the prediction step, while the lower-frequency aFCD speeds guide the slower-varying trend through consistent correction in the update step. Subsection 5.1 details the computation of $\mathcal{K}_{t}(\theta)$, and the adaptations to enforce spatial transferability are described in Subsection 5.2.